OpenStudy (anonymous):

(3a^3 − 5b^3) +_____ a^3 +_____ b^3 = (a^3 + b^3). Type the correct answer in each box. Use numerals instead of words. If necessary, use / for the fraction bar(s).

OpenStudy (igreen):

3 + ? = 0 -5 + ? = 0

OpenStudy (anonymous):

-3 & 5..

OpenStudy (igreen):

Yes, those are your answers.

OpenStudy (igreen):

-3a^3 + 5b^3

OpenStudy (igreen):

Oops, my bad..lol. It kshould be equal to 1, not zero.

OpenStudy (anonymous):

;D Thanks

OpenStudy (igreen):

So it's 3 + ? = 1

OpenStudy (igreen):

-5 + ? = 1

OpenStudy (igreen):

xD My bad

OpenStudy (anonymous):

-2, 6

OpenStudy (igreen):

Yes, those are our answers..lol. -2a^3 + 6b^3

OpenStudy (igreen):

3a^3 - 2a^3 = a^3 -5b^3 + 6b^3 = b^3

OpenStudy (igreen):

You're welcome, and you shouldn't make "Qualified Helper" questions if there aren't any online.

OpenStudy (igreen):

@Preetha doesn't really help..so you should wait for someone else to get on.

OpenStudy (anonymous):

well I figured he was one and would have helped but now I know, why aren't you one?

OpenStudy (igreen):

she* I don't know, I applied..I guess they don't want me. Eh, whatever :P